# Designing Adaptive User Interfaces for mHealth applications targeting chronic disease: A User-Centric Approach

### v1

@@ 7. Threats to Validity
External validity. For the interview and focus group study, we employed STGT’s theoretical sampling technique, which is aimed at generating insights from representative samples rather than achieving generalisability (Hargittai and Micheli, 2019; Hoda, 2021). Therefore, the qualitative findings of our study cannot be generalised to broader populations, but offer valuable insights into how users perceive AUIs in mHealth applications for individuals with chronic diseases. Future research endeavours should seek to validate our findings with more diverse populations, particularly those with lower levels of digital literacy (Hargittai and Micheli, 2019). Secondly, the use of a high-fidelity design prototype in our user study may have inadvertently introduced discrepancies between the prototyped adaptations and the practical applications. This prototype may not fully capture the intricacies of real-world interactions, such as the dynamic interaction between adaptations and other system functionalities or fail to account for users’ nuanced allocation of time to specific adaptations (Bunt et al., 2004; Gajos et al., 2006, 2008). These complexities are crucial to consider as they could influence users’ perceptions and behaviours in ways that the prototype might not accurately reflect. Therefore, future research efforts should explore ways to bridge this gap between high-fidelity prototypes and real-world implementations to ensure that the findings are applicable and actionable in practical settings. Thirdly, we observed demographic differences in our study concerning nationality, age, and the clinic-based population. Although we shared the survey on social media to obtain worldwide participation, we could not achieve it and found that the majority of the participants (49%) were from Australia (Table 4). Hence, similar to the study (Robbins et al., 2017), our findings may not be generalisable to the global population of individuals with chronic diseases. Our survey sample exhibited a skew towards younger demographics compared to the participants involved in the interview and focus group studies. This demographic disparity is significant as patterns of usage and preference are likely to vary across different age groups, with older populations potentially exhibiting different behaviours and preferences. Additionally, it is important to acknowledge that our study surveyed a general population rather than specifically targeting individuals known to be in hospital or clinic-based settings. These individuals, who regularly engage with healthcare services, may exhibit different behaviours and usage patterns of mHealth applications compared to those of the general population with milder symptoms.
@@
Internal validity. Within our prototype, we curated a diverse range of adaptations. However, participants may have encountered instances of confusion when interacting with certain elements of the AUIs. To mitigate this issue, we incorporated an explanatory video to aid participants in navigating through the various adaptations. Furthermore, we supplemented this with detailed instructions for reference, ensuring participants had guidance readily available should they encounter any challenges while using the system. Similarly, we anticipated potential complexities during the survey phase, and thus provided two concrete examples of AUIs to clarify the AUI concepts for participants. We implemented the STGT approach for qualitative data analysis and facilitated extensive team discussions to review and refine our analyses, findings, and presentation, thus mitigating potential biases. Although offering compensation to participants for their participation in interviews and focus group studies can raise concerns about the potential of participants to provide false information to qualify for the study (Head, 2009), it is noteworthy that four participants declined compensation. Instead, they expressed a preference for the funds to be allocated towards further research endeavours. This underscores the voluntary nature of our study and emphasises the perceived importance of the research from the perspective of participants. The survey responses regarding opinions on adaptations for mHealth apps may shift based on their evolving needs and experiences as users interact with mHealth apps over time, and they may also stop using the app, as long-term use of the mHealth app has always been an issue (Woldaregay et al., 2018).
Conclusion validity. While our study focuses on various chronic diseases, it is important to note that the findings may have limitations in terms of generalisability. Nevertheless, our approach aligns with the common practice of including individuals with multiple chronic diseases in research (Elzen et al., 2007). In fact, some participants in our study presented with multiple chronic diseases, mirroring real-life scenarios where individuals often contend with comorbidity alongside their primary condition (di San Marco et al., 2019; Islam et al., 2014). Moving forward, the practical utility of our guidelines warrants validation through user evaluations of adaptive mHealth apps. Our study represents a preliminary evaluation of the framework, and a larger-scale follow-up study involving industry practitioners could provide valuable insights. It is essential to recognise that our framework is evolutionary in nature, and we encourage further research to explore its applicability to population groups at the intersections and with unique needs.
@@
In response to society’s demand for improved mHealth applications to assist diverse users in managing chronic diseases, our study explored user perspectives on various Adaptive User Interface (AUI) approaches in chronic disease management apps. Through this research, we identified significant challenges associated with the use of AUIs and proposed strategies to enhance adaptation design through interviews and focus group studies. Our analysis was further supported by a post-session survey, which corroborated the emergence of these categories. However, it is essential to acknowledge that the adoption of these strategies may entail certain costs, influenced by variables such as user involvement levels, prior experience with mHealth apps, and individual health conditions. Furthermore, we conducted an in-depth exploration of the relationship between specific challenges and the corresponding strategies proposed, illuminating the nuanced trade-offs involved. This provided valuable insights into prioritising user and app characteristics for optimal adaptation design. The quantitative survey provided additional key information on emerging categories and revealed participants’ preferences for different adaptations, data types, collection methods, and levels of engagement. Finally, these insights and categories were synthesised and mapped to existing guidelines or design considerations for mHealth app design adaptation to distil a set of guidelines. We refined these guidelines with 20 participants. These guidelines, which serve as a roadmap for the design of future AUIs in mHealth apps targeting chronic disease development, align with existing design principles, ensuring their relevance and applicability in real-world settings. We characterise this effort in refining guidelines as a preliminary evaluation, and a comprehensive follow-up investigation involving professionals from the industry and academic could be the subsequent phase. It is important to note that our framework is progressive and we advocate for additional research to expand or examine its applicability to different demographic groups.
@@ Acknowledgements.
Wang, Grundy and Madugalla are supported by ARC Laureate Fellowship FL190100035. We would like to acknowledge Daniel Gaspar Figueiredo, Elton Lobo, Michael Wheeler, and Paul Jansons for helping us recruit participants for the user study. Special thanks to all the participants for sharing their valuable perspectives and experiences. The following colleagues provided valuable assistance in the form of comments on earlier drafts, data analysis, prototype evaluation and/or graphics in the study: Dulaji Hidellaarachchi, Elizabeth Manias, Kashumi Madampe and Weimin Wang. We are grateful to Rashina Hoad for her assistance in applying Socio-Technical Grounded Theory (STGT) for data analysis in our paper.References
@@
- Akbar et al. (2020) Saba Akbar, Enrico Coiera, and Farah Magrabi. 2020. Safety concerns with consumer-facing mobile health applications and their consequences: a scoping review. Journal of the American Medical Informatics Association 27, 2 (2020), 330–340.
@@
- Albahri et al. (2018) Ahmed Shihab Albahri, AA Zaidan, Osamah Shihab Albahri, BB Zaidan, and MA Alsalem. 2018. Real-time fault-tolerant mHealth system: Comprehensive review of healthcare services, opens issues, challenges and methodological aspects. Journal of medical systems 42 (2018), 1–56.
@@
- Alvarez-Cortes et al. (2007) Victor Alvarez-Cortes, Benjamin E Zayas-Perez, Victor Huga Zarate-Silva, and Jorge A Ramirez Uresti. 2007. Current trends in adaptive user interfaces: Challenges and applications. In Electronics, Robotics and Automotive Mechanics Conference (CERMA 2007). IEEE, 312–317.
@@
- Attfield et al. (2011) S Attfield, G Kazai, M Lalmas, and B Piwowarski. 2011. Towards a science of user engagement. WSDM workshop on user modelling for web applications.
@@
- Baroni et al. (2022) Ilaria Baroni, Gloria Re Calegari, Damiano Scandolari, and Irene Celino. 2022. AI-TAM: a model to investigate user acceptance and collaborative intention in human-in-the-loop AI applications. Human Computation 9, 1 (2022), 1–21.
@@
- Bentley et al. (2014) Claire L Bentley, Lauren A Powell, Alison Orrell, and Gail A Mountain. 2014. Addressing design and suitability barriers to telecare use: has anything changed? Technology and Disability 26, 4 (2014), 221–235.
- Bettman (1971) James R Bettman. 1971. The structure of consumer choice processes. Journal of marketing research 8, 4 (1971), 465–471.
@@
- Cajita et al. (2018) Maan Isabella Cajita, Nancy A Hodgson, Katherine Wai Lam, Sera Yoo, and Hae-Ra Han. 2018. Facilitators of and barriers to mHealth adoption in older adults with heart failure. Computers, informatics, nursing: CIN 36, 8 (2018), 376.
@@
- Cheverst et al. (2005) Keith Cheverst, Hee Eon Byun, Dan Fitton, Corina Sas, Chris Kray, and Nicolas Villar. 2005. Exploring issues of user model transparency and proactive behaviour in an office environment control system. User Modeling and User-Adapted Interaction 15 (2005), 235–273.
@@
- Codish and Ravid (2014) David Codish and Gilad Ravid. 2014. Personality based gamification-Educational gamification for extroverts and introverts. In Proceedings of the 9th CHAIS Conference for the Study of Innovation and Learning Technologies: Learning in the Technological Era, Vol. 1. The Open University of Israel Ra’anana, 36–44.
@@
- Findlater and McGrenere (2010) Leah Findlater and Joanna McGrenere. 2010. Beyond performance: Feature awareness in personalized interfaces. Int. J. Hum. Comput. Stud. 68 (2010), 121–137. https://api.semanticscholar.org/CorpusID:10129897
@@
- Folkman and Lazarus (1980) Susan Folkman and Richard S Lazarus. 1980. An analysis of coping in a middle-aged community sample. Journal of health and social behavior (1980), 219–239.
@@
- Glaser (1992) Barney G Glaser. 1992. Basics of grounded theory analysis: Emergence vs forcing. (No Title) (1992).
@@
- Haggag et al. (2022) Omar Haggag, John Grundy, Mohamed Abdelrazek, and Sherif Haggag. 2022. A large scale analysis of mHealth app user reviews. Empirical Software Engineering 27, 7 (2022), 196.
@@
- Hammond (1987) Nick Hammond. 1987. Principles from the psychology of skill acquisition. In Applying cognitive psychology to user-interface design. 163–187.
@@
- Kelders et al. (2012) Saskia M Kelders, Robin N Kok, Hans C Ossebaard, and Julia EWC Van Gemert-Pijnen. 2012. Persuasive system design does matter: a systematic review of adherence to web-based interventions. Journal of medical Internet research 14, 6 (2012), e152.
@@
- Kurtz (2005) Matthew M Kurtz. 2005. Neurocognitive impairment across the lifespan in schizophrenia: an update. Schizophrenia research 74, 1 (2005), 15–26.
@@
- Li et al. (2020) Jiaoyang Li, Cheng Zhang, Xixi Li, and Chenghong Zhang. 2020. Patients’ emotional bonding with MHealth apps: An attachment perspective on patients’ use of MHealth applications. International Journal of Information Management 51 (2020), 102054.
@@
- Lohr (2015) Steve Lohr. 2015. The healing power of your own medical records. The New York Times 31, 3 (2015), 1–8.
@@
- McDougall et al. (2000) Siné JP McDougall, Oscar De Bruijn, and Martin B Curry. 2000. Exploring the effects of icon characteristics on user performance: The role of icon concreteness, complexity, and distinctiveness. Journal of Experimental Psychology: Applied 6, 4 (2000), 291.
@@
- Obie et al. (2021) Humphrey O Obie, Mojtaba Shahin, John Grundy, Burak Turhan, Li Li, Waqar Hussain, and Jon Whittle. 2021. Does domain change the opinion of individuals on human values? A preliminary investigation on eHealth apps end-users. In 2021 28th Asia-Pacific Software Engineering Conference (APSEC). IEEE, 531–535.
@@
- Oertel et al. (2020) Catharine Oertel, Ginevra Castellano, Mohamed Chetouani, Jauwairia Nasir, Mohammad Obaid, Catherine Pelachaud, and Christopher Peters. 2020. Engagement in human-agent interaction: An overview. Frontiers in Robotics and AI 7 (2020), 92.
@@
- Parker et al. (2019) Charlie Parker, Sam Scott, and Alistair Geddes. 2019. Snowball sampling. SAGE research methods foundations (2019).
@@
- Philip et al. (2022) Ben Joseph Philip, Mohamed Abdelrazek, Alessio Bonti, Scott Barnett, and John Grundy. 2022. Data collection mechanisms in health and wellness apps: review and analysis. JMIR mHealth and uHealth 10, 3 (2022), e30468.
@@
- Prasad et al. (2014) Aarathi Prasad, Jacob Sorber, Timothy Stablein, D Anthony, and David Kotz. 2014. Understanding user privacy preferences for mhealth data sharing. mHealth multidisciplinary verticals 30 (2014), 545–569.
@@
- Ramos et al. (2019) Gonzalo Ramos, Jina Suh, Soroush Ghorashi, Christopher Meek, Richard Banks, Saleema Amershi, Rebecca Fiebrink, Alison Smith-Renner, and Gagan Bansal. 2019. Emerging perspectives in human-centered machine learning. In Extended Abstracts of the 2019 CHI Conference on Human Factors in Computing Systems. 1–8.
@@
- Sara (2021) Berg Sara. 2021. What doctors wish patients knew about decision fatigue. https://www.ama-assn.org/delivering-care/public-health/what-doctors-wish-patients-knew-about-decision-fatigue

### v2

@@ 7.1. External Validity
Stage One. For participant recruitment in the interview and focus group study, the Socio-Technical Grounded Theory (STGT) theoretical sampling technique was employed. This approach is iterative and supports interleaved data collection and analysis (Hargittai and Micheli, 2019; Hoda, 2021). As a result, the qualitative findings of this study are not intended to be generalizable to larger populations. Instead, they provide in-depth insight into how individuals with chronic diseases perceive and interact with Adaptive User Interfaces (AUIs) in mHealth applications. Future research efforts should seek to validate our findings with more diverse populations, particularly those with lower levels of digital literacy (Hargittai and Micheli, 2019). Second, the use of a prototype in stage one may have inadvertently introduced discrepancies between the prototyped adaptations and the practical applications. This prototype may not fully capture the intricacies of real-world interactions, such as the dynamic interaction between adaptations and other system functionalities or fail to account for users’ nuanced allocation of time to specific adaptations (Bunt et al., 2004; Gajos et al., 2006, 2008). These complexities are crucial to consider, as they could influence users’ perceptions and behaviors in ways that the prototype might not accurately reflect. Third, we observed demographic differences in our user survey in stage one study with respect to nationality, age, and the clinic population. Although we shared the survey on social media to obtain worldwide participation, we were unable to achieve it and found that the majority of the participants (49%) were from Australia (Table 1). Hence, similar to the observations in (Robbins et al., 2017), the findings of this study may have limited generalizability beyond the specific group of participants. Our user survey sample exhibited a skew towards younger demographics. This demographic disparity is significant as usage and preference patterns are likely to vary between different age groups, and older populations can potentially exhibit different behaviors and preferences. Furthermore, it is important to acknowledge that our study surveyed a general population rather than individuals in hospital or clinic settings. These individuals, who regularly engage with healthcare services, may exhibit different behaviors and usage patterns of mHealth applications compared to those of the general population with milder symptoms.
Stage Two. The guideline evaluation study raises concerns about generalizability due to diverse practitioner backgrounds, organizational practices, and target user groups. There are several limitations identified in this case study. First, the limited number of evaluators, only two evaluators, constrained the study, despite recommendations to use at least four evaluators for identifying most issues (Nielsen, 1995). Second, while both evaluators had experience in UX/UI design and front-end development, they may not represent a broader sample of nonexpert users, limiting the generalizability of the evaluation. Third, evaluating only four diabetes-related mHealth applications restricts the extent of insights gained, as a comprehensive evaluation encompassing a wider spectrum of available applications could more effectively evaluate the efficacy of the guidelines. In addition, the diverse nature of chronic diseases and user demographics complicates the creation of universal design guidelines. The guidelines aim to be flexible and cover common needs across chronic diseases, but might lack specificity when applied to particular diseases or specific user groups. We recommend that future studies address these limitations by including a larger and more diverse group of evaluators and testing the guidelines in a broader range of usage contexts and devices to improve their applicability and robustness. Although it recognizes the importance of understanding the underlying purpose of guidelines for individuals with chronic diseases, a detailed investigation of this topic lies beyond the scope of this work. Future research should explore how different user-related factors, especially in individuals with chronic diseases, impact the design of adaptations.
7.2. Internal Validity
@@
Stage One. Within the AUI prototype, some participants experienced confusion when interacting with specific adaptive elements. To address this, an explanatory video was included to guide users and clarify the purpose and functionality of the different adaptations. Furthermore, we supplemented this with detailed instructions for reference, ensuring that participants had readily available guidance should they encounter any challenges while using the system. Similarly, we anticipated potential complexities during the survey phase and thus provided two concrete examples of AUIs to clarify AUI concepts for participants. We used the STGT approach for qualitative data analysis and facilitated extensive team discussions to review and refine our analyses, findings, and presentation, thus mitigating potential biases. Although offering compensation to participants for their participation in interviews and focus group studies can raise concerns about the potential of participants to provide false information to qualify for the study (Head, 2009), it should be noted that four participants declined compensation. Instead, they expressed a preference for the funds to be allocated towards further research endeavors. This underscores the voluntary nature of our study. The survey responses regarding opinions on adaptations for mHealth applications may change based on their evolving needs and experiences as users interact with mHealth applications over time, and they may also stop using them, as long-term use of the mHealth application has always been an issue (Woldaregay et al., 2018).
Stage Two. Software practitioners with different levels of experience in accessibility or adaptive design can provide inconsistent feedback, and personal preferences or biases toward specific design approaches could influence their evaluations. Furthermore, the qualitative nature of the study introduces subjectivity, as interpretations of the guidelines may vary among software practitioners. To mitigate this limitation, the guideline evaluation survey included two comprehension check questions designed to assess participants’ understanding of the guidelines. Practitioners who answered incorrectly were given the opportunity to review the question and the guidelines before proceeding, helping to ensure more informed and consistent responses.
7.3. Conclusion Validity
We acknowledge the limitation of using a single prototype to address a wide range of chronic diseases, as different conditions often entail varied user needs, interaction requirements, and adaptation preferences. This choice may constrain the generalizability of the findings. However, our approach is consistent with the common practice of including individuals with multiple chronic diseases in research (Elzen et al., 2007). In fact, some participants in our study presented multiple chronic diseases, mirroring real-life scenarios in which people often contend with comorbidity alongside their primary disease (di San Marco et al., 2019; Islam et al., 2014). Involving the same participants in multiple stages of the study, interviews, focus groups, and end-user guideline evaluations can introduce bias due to their previous exposure to the study context. This repeated participation could influence responses and evaluations, potentially reducing confidence in the generalization and robustness of the results.
@@ 8. Conclusion
This study addressed the pressing need for more user-centered mHealth applications to support individuals managing chronic diseases. Conducted in two stages, the research began with a user study, comprising interviews, focus groups, and surveys, to examine how users perceive and interact with Adaptive User Interfaces (AUIs) in mHealth contexts. Insights from this stage informed the development of a new set of guidelines, which were further refined through feedback from end-users and software practitioners. These guidelines were subsequently validated through case studies involving four real-world mHealth applications, with additional analysis of user reviews to determine alignment between expert-identified issues and end-user experiences. The results suggest that the proposed guidelines are more effective in uncovering critical usability and adaptation challenges than the existing mHealth usability guidelines. The nine guidelines, organized into four groups, offer software practitioners a structured framework to design adaptive features that accommodate user variability while supporting long-term use. Several avenues remain for further exploration. One potential avenue is the development of automated tools, design templates, or UI component libraries aligned with the guidelines could support their practical integration into development workflows. Another is conducting longitudinal field studies or randomized controlled trials that could evaluate the long-term impact of guideline-based AUI implementations on user engagement, health outcomes, and system usability.
@@ Acknowledgements.
Wang, Grundy and Madugalla are supported by the ARC Laureate Fellowship FL190100035. We would like to acknowledge Daniel Gaspar Figueiredo, Elton Lobo, Michael Wheeler, and Paul Jansons for helping us recruit participants for the user study. We extend our sincere thanks to all the end-user participants for sharing their valuable experiences and perspectives, and to the software practitioners who contributed insightful feedback that helped strengthen the proposed guidelines. Special thanks to Rashina Hoda for her guidance in applying Socio-Technical Grounded Theory (STGT) for data analysis and to Md Shamsujjoha for helping refine the presentation of the guideline section in this paper. The following colleagues provided valuable assistance in the form of comments on earlier drafts, data analysis, prototype evaluation, and/or graphics in the study: Dulaji Hidellaarachchi, Elizabeth Manias, Kashumi Madampe, Suyu Ma and Weimin Wang.References
@@
- Alshayban et al. (2020) Abdulaziz Alshayban, Iftekhar Ahmed, and Sam Malek. 2020. Accessibility issues in android apps: state of affairs, sentiments, and ways forward. In Proceedings of the ACM/IEEE 42nd International Conference on Software Engineering. 1323–1334.
- Alsswey and Al-Samarraie (2021) Ahmed Alsswey and Hosam Al-Samarraie. 2021. The role of Hofstede’s cultural dimensions in the design of user interface: the case of Arabic. Ai Edam 35, 1 (2021), 116–127.
@@
- Athilingam et al. (2018) Ponrathi Athilingam, Bradlee Jenkins, et al. 2018. Mobile phone apps to support heart failure self-care management: integrative review. JMIR cardio 2, 1 (2018), e10057.
@@
- Board (2019) Xcertia Board. 2019. Xcertia - mHealth app guideline. Retrieved December 28, 2024 from https://www.himss.org/sites/hde/files/media/file/2020/04/17/xcertia-guidelines-2019-final.pdf
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using thematic analysis in psychology. Qualitative research in psychology 3, 2 (2006), 77–101.
- Braun and Clarke (2012) Virginia Braun and Victoria Clarke. 2012. Thematic analysis. American Psychological Association.
@@
- Burgess et al. (2019) Eleanor R Burgess, Madhu C Reddy, Andrew Davenport, Paul Laboi, and Ann Blandford. 2019. ” Tricky to get your head around” Information Work of People Managing Chronic Kidney Disease in the UK. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. 1–17.
@@
- Campbell (2015) Ollie Campbell. 2015. Designing for the elderly: Ways older people use digital technology differently. Retrieved February 28, 2025 from https://www.smashingmagazine.com/2015/02/designing-digital-technology-for-the-elderly/
@@
- Castelli et al. (2017) Nico Castelli, Corinna Ogonowski, Timo Jakobi, Martin Stein, Gunnar Stevens, and Volker Wulf. 2017. What happened in my home? an end-user development approach for smart home data visualization. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems. 853–866.
- Chen et al. (2024) Mengyu Chen, Shuang He, and Xinyi Fu. 2024. A Case Study Exploring the Applicability of Heuristic Evaluation in Smart Home Systems. In Extended Abstracts of the CHI Conference on Human Factors in Computing Systems. 1–7.
- Chib and Lin (2018) Arul Chib and Sapphire H Lin. 2018. Theoretical advancements in mHealth: a systematic review of mobile apps. Journal of health communication 23, 10-11 (2018), 909–955.
- Choe et al. (2017) Eun Kyoung Choe, Saeed Abdullah, Mashfiqui Rabbi, Edison Thomaz, Daniel A Epstein, Felicia Cordeiro, Matthew Kay, Gregory D Abowd, Tanzeem Choudhury, James Fogarty, et al. 2017. Semi-automated tracking: a balanced approach for self-monitoring applications. IEEE Pervasive Computing 16, 1 (2017), 74–84.
@@
- Conor et al. (2014) Hanrahan Conor, Aungst Timothy, and Cole Sabrina. 2014. Evaluating mobile medical applications. Retrieved December 22, 2024 from https://www.ashp.org/-/media/store%20files/mobile-medical-apps.pdf
@@
[…]
- Daihua et al. (2015) X Yu Daihua, Bambang Parmanto, Brad E Dicianno, and Gede Pramana. 2015. Accessibility of mHealth self-care apps for individuals with spina bifida. Perspectives in health information management 12, Spring (2015).
@@
- Depuy et al. (2023) J Depuy, J Molina-Morales, AJ Conover, and GJ Jerome. 2023. Accessible weight loss program for adults who are legally blind: A pilot study. Disability and Health Journal 16, 4 (2023), 101502.
@@
- Epstein et al. (2016) Daniel A Epstein, Monica Caraway, Chuck Johnston, An Ping, James Fogarty, and Sean A Munson. 2016. Beyond abandonment to next steps: understanding and designing for life after personal informatics tool use. In Proceedings of the 2016 CHI conference on human factors in computing systems. 1109–1113.
@@
- Firmenich et al. (2019) Sergio Firmenich, Alejandra Garrido, Fabio Paternò, and Gustavo Rossi. 2019. User interface adaptation for accessibility. Web accessibility: A foundation for research (2019), 547–568.
@@
- Gandapur et al. (2016) Yousuf Gandapur, Sina Kianoush, Heval M Kelli, Satish Misra, Bruno Urrea, Michael J Blaha, Garth Graham, Francoise A Marvel, and Seth S Martin. 2016. The role of mHealth for improving medication adherence in patients with cardiovascular disease: a systematic review. European Heart Journal–Quality of Care and Clinical Outcomes 2, 4 (2016), 237–244.
- Glaser (1992) Barney G Glaser. 1992. Basics of grounded theory analysis: Emergence vs forcing. Sociology Press (1992).
@@
- Goundar et al. (2024) Munil Shiva Goundar, Bimal Aklesh Kumar, and ABM Shawkat Ali. 2024. Development of usability guidelines: a systematic literature review. International Journal of Human–Computer Interaction 40, 5 (2024), 1298–1316.
@@
- Hoda (2024) Rashina Hoda. 2024. Qualitative Research with Socio-Technical Grounded Theory. SpringerLink.
- Hofstede (2001) Geert Hofstede. 2001. Culture’s consequences: Comparing values, behaviors, institutions, and organizations across nations. International Educational and Professional (2001).
@@
- Horvitz (1999) Eric Horvitz. 1999. Principles of mixed-initiative user interfaces. In Proceedings of the SIGCHI conference on Human Factors in Computing Systems. 159–166.
- Huh and Ackerman (2012) Jina Huh and Mark S Ackerman. 2012. Collaborative help in chronic disease management: supporting individualized problems. In Proceedings of the ACM 2012 conference on Computer Supported Cooperative Work. 853–862.
@@
- Johnsson and Magnusson (2020) Björn A Johnsson and Boris Magnusson. 2020. Towards end-user development of graphical user interfaces for internet of things. Future Generation Computer Systems 107 (2020), 670–680.
@@
- Kim (2019) Hyung Nam Kim. 2019. Mobile health technology accessible to people with visual impairments. Journal on technology and persons with disabilities 7 (2019).
@@
- Langevin et al. (2021) Raina Langevin, Ross J Lordon, Thi Avrahami, Benjamin R Cowan, Tad Hirsch, and Gary Hsieh. 2021. Heuristic evaluation of conversational agents. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. 1–15.
@@
- Mack et al. (2021) Kelly Mack, Emma McDonnell, Dhruv Jain, Lucy Lu Wang, Jon E. Froehlich, and Leah Findlater. 2021. What do we mean by “accessibility research”? A literature survey of accessibility papers in CHI and ASSETS from 1994 to 2019. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. 1–18.
- Mack et al. (2022) Kelly Mack, Emma J McDonnell, Leah Findlater, and Heather D Evans. 2022. Chronically Under-Addressed: Considerations for HCI Accessibility Practice with Chronically Ill People. In Proceedings of the 24th International ACM SIGACCESS Conference on Computers and Accessibility. 1–15.
@@
- Matthew-Maich et al. (2016) Nancy Matthew-Maich, Lauren Harris, Jenny Ploeg, Maureen Markle-Reid, Ruta Valaitis, Sarah Ibrahim, Amiram Gafni, Sandra Isaacs, et al. 2016. Designing, implementing, and evaluating mobile health technologies for managing chronic conditions in older adults: a scoping review. JMIR mHealth and uHealth 4, 2 (2016), e5127.
@@
- Melchior et al. (2012) Jérémie Melchior, Jean Vanderdonckt, and Peter Van Roy. 2012. A comparative evaluation of user preferences for extra-user interfaces. International Journal of Human-Computer Interaction 28, 11 (2012), 760–767.
@@
- Mezhoudi and Vanderdonckt (2021) Nesrine Mezhoudi and Jean Vanderdonckt. 2021. Toward a task-driven intelligent GUI adaptation by mixed-initiative. International Journal of Human–Computer Interaction 37, 5 (2021), 445–458.
- Milne et al. (2014) Lauren R Milne, Cynthia L Bennett, and Richard E Ladner. 2014. The accessibility of mobile health sensors for blind users. In International technology and persons with disabilities conference scientific/research proceedings (CSUN 2014). 166–175.
- Miniukovich et al. (2019) Aliaksei Miniukovich, Michele Scaltritti, Simone Sulpizio, and Antonella De Angeli. 2019. Guideline-based evaluation of web readability. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. 1–12.
@@
- Morey et al. (2017) Stephanie A Morey, Laura H Barg-Walkow, and Wendy A Rogers. 2017. Managing heart failure on the go: usability issues with mHealth apps for older adults. In Proceedings of the Human Factors and Ergonomics Society Annual Meeting, Vol. 61. SAGE Publications Sage CA: Los Angeles, CA, 1–5.
@@
- Nedopil et al. (2013) Ch Nedopil, C Schauber, and S Glende. 2013. Knowledge base: AAL stakeholders and their requirements. Ambient Assisted Living Association: Brussels, Belgium (2013).
@@
- Nielsen (1994) Jakob Nielsen. 1994. Usability inspection methods. In Conference companion on Human factors in computing systems. 413–414.
- Nielsen (1995) Jakob Nielsen. 1995. How to conduct a heuristic evaluation. Nielsen Norman Group 1, 1 (1995), 8.
@@
- NLTK (2024) NLTK. 2024. Natural Language Toolkit. https://www.nltk.org/
@@
- Nunes et al. (2015) Francisco Nunes, Nervo Verdezoto, Geraldine Fitzpatrick, Morten Kyng, Erik Grönvall, and Cristiano Storni. 2015. Self-care technologies in HCI: Trends, tensions, and opportunities. ACM Transactions on Computer-Human Interaction (TOCHI) 22, 6 (2015), 1–45.
@@
- Park and Lee (2015) Kibeom Park and Seok-Won Lee. 2015. Model-based approach for engineering adaptive user interface requirements. In Requirements Engineering in the Big Data Era: Second Asia Pacific Symposium, APRES 2015, Wuhan, China, October 18–20, 2015, Proceedings. Springer, 18–32.
@@
- Paymal and Homewood (2024) Léa Paymal and Sarah Homewood. 2024. Good Days, Bad Days: Understanding the Trajectories of Technology Use During Chronic Fatigue Syndrome. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems. 1–10.
@@
- Radcliffe et al. (2021) Erin Radcliffe, Ben Lippincott, Raeda Anderson, and Mike Jones. 2021. A pilot evaluation of mHealth app accessibility for three top-rated weight management apps by people with disabilities. International Journal of Environmental Research and Public Health 18, 7 (2021), 3669.
@@
- Reyes Arias et al. (2022) Jose E Reyes Arias, Kale Kurtzhall, Di Pham, Mohamed Wiem Mkaouer, and Yasmine N Elglaly. 2022. Accessibility feedback in mobile application reviews: A dataset of reviews and accessibility guidelines. In CHI Conference on Human Factors in Computing Systems Extended Abstracts. 1–7.
@@
- Ross et al. (2017) Anne Spencer Ross, Xiaoyi Zhang, James Fogarty, and Jacob O Wobbrock. 2017. Epidemiology as a framework for large-scale mobile application accessibility assessment. In Proceedings of the 19th international ACM SIGACCESS conference on computers and accessibility. 2–11.
- Ryu et al. (2023) Hyeyoung Ryu, Andrew BL Berry, Catherine Y Lim, Andrea Hartzler, Tad Hirsch, Juanita I Trejo, Zoë Abigail Bermet, Brandi Crawford-Gallagher, Vi Tran, Dawn Ferguson, et al. 2023. “You Can See the Connections”: Facilitating Visualization of Care Priorities in People Living with Multiple Chronic Health Conditions. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–17.
@@
- Schobel et al. (2016) Johannes Schobel, Rüdiger Pryss, Marc Schickler, Martina Ruf-Leuschner, Thomas Elbert, and Manfred Reichert. 2016. End-user programming of mobile services: empowering domain experts to implement mobile data collection applications. In 2016 IEEE International Conference on Mobile Services (MS). IEEE, 1–8.
- Sefidgar et al. (2024) Yasaman S Sefidgar, Carla L Castillo, Shaan Chopra, Liwei Jiang, Tae Jones, Anant Mittal, Hyeyoung Ryu, Jessica Schroeder, Allison Cole, Natalia Murinova, et al. 2024. MigraineTracker: Examining Patient Experiences with Goal-Directed Self-Tracking for a Chronic Health Condition. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–19.